# Towards AI-Architecture Liberty: A Comprehensive Survey on Designing and Collaborating Virtual Architecture by Deep Learning in the Metaverse

### v3

@@ 5.3. Enhancing User Consideration with Diversity and Collectives under Designer-AI Collaboration
Designs involving multiple users necessitate inclusivity by incorporating usage affordances and considering wide needs for virtual architecture since the metaverse pertains to common environments for collectives and individuals from diverse backgrounds and cultures through converging and engaging, leading to highly multi-social experiences (Jonassen and Rohrer-Murphy, 1999; Bartle, 2003). Accordingly, we identify two focal conflicts regarding this agenda: the balance between collective and individual, as well as culture and efficiency.
@@
Individuals and Collectives. A significant challenge arises from the conflict between individual-centric ideologies of new liberalism and the collective nature of the metaverse (Nazmeeva, 2019). While the exhibiting of ownership, identity, and recommended content prominence the precise value of an individual in expansive virtual spaces under new liberalism (Nazmeeva, 2019). Within this dynamic environment, we are increasingly accused of proliferating products manifesting our unique identities and personal needs (Nazmeeva, 2019). This situation leads to an urgent rethinking of the collaboration with virtual worlds regarding the relationship between collectivism and individualism.
One potential solution is public participation in the design process. Designers have conducted participatory design experiments positively, addressing multifaceted problems and designing for minority groups. Participatory design allows non-professionals to make design decisions collaboratively alongside professional designers. The emergence of creative collectives in architecture underscores the importance of trends in public participation (Eloy et al., 2021). However, our analysis of the survey found no research on participatory concerns for 3D generation or DL-assisted architectural design. In contrast, there has been a significant achievement in participation topics within human-AI collaboration (Simonsen et al., 2020). Therefore, we urge designers and researchers to emphasize the public participatory potential for virtual building by referencing existing research.
5.4. Designing Adaptive Tool by Integrating Multiple Tools with Multimodal User Input and Output.
Integrating tools results in wide access with a user-friendly interface towards rich user-created content in the future metaverse. Although this vision has been underscored in the metaverse industry, the adaptive tool for diverse users to design fields is unseasoned. The reason behind this agenda is the lack of applications that integrate multimodality and consider user experience.
@@ 5.5. Research Directions and Future Opportunities
5.5.1. Designing User-Centered Adaptive Virtual Buildings
@@
Virtual buildings, serving as 3D digital resources of content and space, play a pivotal role in fostering the metaverse content ecosystem, which involves consumers, creators, and platforms. In a recent study, Ding et al. exhibited an innovative approach to user-personalized architectural design in the virtual space, utilizing a user-friendly tool for human-AI collaboration (Ding et al., 2022b). This tool showcases an interface and features tailored for creating virtual architecture, aiming to ensure accessibility and ease of use. Tools featuring user-friendly interfaces and customization options can empower consumers to assume governance and content-generation roles. Therefore, we believe multiple approaches enabling user engagement in various roles represent the prevailing trend within the metaverse context.
@@
Deep learning is propelling virtual architecture towards innovative forms shaped by algorithmic principles, surpassing conventional visual geometries. The algorithmic form encapsulates the relationship between computational processes and the resulting object characteristics (Chaitin, 1975b, a). For instance, recent research has proliferated in aesthetics assessment (Tonn, 2017; Kavakoglu, 2021; Sardenberg and Theron Burger, 2019), regarding visual character (Stuart-Smith and Danahy, 2022), and visual sense (Wells et al., 2021). In this context, an increasing number of social algorithm proposals integrate social factors like environmental and social sustainability into deep learning frameworks. For instance, Koh et al. employed 3D GANs to design a series of housing buildings ensuring sustainable economic balance and social resident privacy (Koh, 2022). Moreover, the instantaneous feedback loop in environmentally intelligent virtual architecture aids its management and governance, streamlining activities like building reviews and further adaptation. This responsive capability fosters a collaborative environment within the metaverse.
@@ 6. Conclusion
In this survey, we explore various approaches deep neural networks utilize to generate virtual architectures autonomously. Our investigation spans three main domains: architectural design, 3D generation techniques, and virtual environments for human-AI collaboration. Through an in-depth literature analysis, we focus on four key aspects: datasets, multimodality, design intuition, and generative frameworks. We aim to bridge the current research gap from an interdisciplinary perspective.
@@
Our survey highlights the lack of systematic research, particularly concerning the virtual dimensionality of architecture and the collaborative design aspects of human-AI interaction. We advocate for comprehensive and interdisciplinary research efforts encompassing various themes such as agency, communication, user considerations, tool integration, and the involvement of diverse stakeholders, including technical developers, HCI researchers, architects, and metaverse practitioners.

### v4

@@ 5.3. Enhancing User Consideration with Diversity and Collectives under Designer-AI Collaboration
Designs involving multiple users necessitate inclusivity by incorporating usage affordances and considering wide needs for virtual architecture since the virtual environment pertains to shared environments for collectives and individuals from diverse backgrounds and cultures through converging and engaging, leading to highly multi-social experiences (Jonassen and Rohrer-Murphy, 1999; Bartle, 2003). Accordingly, we identify two focal conflicts regarding this agenda: the balance between collective and individual, as well as culture and efficiency.
@@
Individuals and Collectives. A significant challenge arises from the conflict between individual-centric ideologies of new liberalism and the collective nature of the virtual environment (Nazmeeva, 2019). While the exhibiting of ownership, identity, and recommended content prominence the precise value of an individual in expansive virtual spaces under new liberalism (Nazmeeva, 2019). Within this dynamic environment, we are increasingly accused of proliferating products manifesting our unique identities and personal needs (Nazmeeva, 2019). This situation leads to an urgent rethinking of the collaboration with virtual worlds regarding the relationship between collectivism and individualism.
One potential solution is public participation in the design process. Designers have conducted participatory design experiments positively, addressing multifaceted problems and designing for minority groups. Participatory design allows non-professionals to make design decisions collaboratively alongside professional designers. The emergence of creative collectives in architecture underscores the importance of trends in public participation (Eloy et al., 2021). However, our analysis of the survey found no research on participatory concerns for 3D generation or DL-assisted architectural design. In contrast, there has been a significant achievement in participation topics within designer-AI collaboration (Simonsen et al., 2020). Therefore, we urge designers and researchers to emphasize the public participatory potential for virtual building by referencing existing research.
5.4. Designing Adaptive Tool by Integrating Multiple Tools with Multimodal User Input and Output
Integrating tools results in wide access with a user-friendly interface towards rich user-created content in the future metaverse. Although this vision has been underscored in the virtual environment industry, the adaptive tool for diverse users to design fields is unseasoned. The reason behind this agenda is the lack of applications that integrate multimodality and consider user experience.
@@ 5.5. Research Directions and Future Opportunities
5.5.1. Designing User-Centric Adaptive Virtual Buildings
@@
Virtual buildings, serving as 3D digital resources of content and space, play a pivotal role in fostering the content ecosystem in the virtual environment, which involves consumers, creators, and platforms. In a recent study, Ding et al. demonstrated an innovative approach to user-personalized architectural design in virtual space, utilizing a user-friendly tool for designer-AI collaboration (Ding et al., 2022b). This tool showcases an interface and features tailored for creating virtual architecture, aiming to ensure accessibility and ease of use. Tools featuring user-friendly interfaces and customization options can empower consumers to assume governance and content-generation roles. Therefore, we believe multiple approaches enabling user engagement in various roles represent the prevailing trend within the virtual environment context.
@@
Deep learning is propelling virtual architecture towards innovative forms shaped by algorithmic principles, surpassing conventional visual geometries. The algorithmic form encapsulates the relationship between computational processes and the resulting object characteristics (Chaitin, 1975b, a). For instance, recent research has proliferated in aesthetics assessment (Tonn, 2017; Kavakoglu, 2021; Sardenberg and Theron Burger, 2019), regarding visual character (Stuart-Smith et al., 2022), and visual sense (Wells et al., 2021). In this context, an increasing number of social algorithm proposals integrate social factors like environmental and social sustainability into deep learning frameworks. For instance, Koh et al. employed 3D GANs to design a series of housing buildings ensuring sustainable economic balance and social resident privacy (Koh, 2022). Moreover, the instantaneous feedback loop in environmentally intelligent virtual architecture aids its management and governance, streamlining activities like building reviews and further adaptation. This responsive capability fosters a collaborative environment.
5.5.4. Ethical implications of AI-generated architecture
In an era of increased collaboration with AI, it is essential to examine the ethical implications of originality and authorship, the morality of machine creation, and human rights within a pervasive data environment. In this context, user privacy and authorship are of utmost importance.
Designers must work closely with AI experts and developers to ensure data privacy, including confidentiality, integrity, and availability. As designers become more involved in AI procedures such as deployment, training, and decision-making as the party of developers, they have a responsibility to raise awareness and take action to protect data privacy (Golda et al., 2024).
AI-generated content in virtual digital environments presents significant challenges related to authorship, ownership, and copyright within a complex content ecosystem due to zero-cost reproduction and machine intelligence (Vyas, 2022). For example, original creations can be utilized by AI models to produce suspected pirated content without informing the authors. While the industry offers solutions such as uploading creations with copyright-level options, challenges persist regarding evaluation and feasibility, particularly in specific areas (Vyas, 2022). Specifically, the reuse of creations should involve informing the author and obtaining permission. The benefits of distribution or authorship should be considered, especially when users may serve in multiple roles, including potential creators, re-creators, distributors, and regulators. Therefore, designers with a robust understanding of the field and a commitment to authorship play a crucial role in establishing specific criteria to regulate the generation approaches of AI-generated architecture.
@@ 6. Conclusion
In this survey, we explore various approaches deep neural networks utilize to generate virtual architectures autonomously. Our investigation spans three main domains: architectural design, 3D generation techniques, and virtual environments for designer-AI collaboration. Through an in-depth literature analysis, we focus on four key aspects: datasets, multimodality, design intuition, and generative frameworks. We aim to bridge the current research gap from an interdisciplinary perspective.
@@
Our survey highlights the lack of systematic research, particularly concerning the virtual dimensionality of architecture and the collaborative design aspects of human-AI interaction. We advocate for comprehensive and interdisciplinary research efforts encompassing various themes such as agency, communication, user considerations, tool integration, and the involvement of diverse stakeholders, including technical developers, HCI researchers, architects, and virtual environment practitioners.
Appendix
Appendix A Supplementary Material (e-pub only)
This material is supplementary and should be included in the electronic version only.
Table 1. Overview of the mission and scope of virtual architecture, focusing on five key aspects: spatial forms, user presence, socialization, interaction, and interoperability and scalability.
Discipline Description
Spatial forms Incorporating diverse forms in terms of function, style, and purpose to ensure efficient production to achieve large-scale outcomes.
User presence, experience Ensuring self-presence with a pleasant, understandable user experience, considering human perception (e.g., comfort, safety), behavioral demands (e.g., navigation, gathering activities), and environmental factors (e.g., lighting, sounds, atmosphere) in immersive environments.
Socialization, engagement Supporting diverse and multiple social interactions, including activities, engagements, and goals for live, entertainment, and work. Providing collaboration opportunities with integrated tools for editing, altering, and customizing tasks.
Interactive elements Designing flexible, dynamic, and real-time interactive human-centric buildings by utilizing ubiquitous data from environmental intelligence, such as responsive or interactive building elements.
Interoperability, scalability Building interoperability and scalability across multiple platforms focusing on data security, user privacy, and resource management for sustainability.
Table 2. This article elucidates four key focuses derived from the research questions reviewed for deep learning-assisted architectural design papers: datasets, multimodality, design intuition, and generative models (depicted as four columns on the right). A selection of the reviewed papers is provided in the table as a reference.
Reference Category Research Question 3D Representation Deep Generative Model Key Focus
Dataset Multi-Modality Design Intuition Generative Model
(Çakmak, 2022) 3D Solid Extend design cognition Point cloud/Mesh GAN (with Autoencoder)
(Mueller, 2023) 3D Solid Generate building massing Point cloud/Voxel Grid GAN
(Miguel et al., 2019) 3D Solid Generation, manipulation and form finding of structural typologies Wireframe VAE
(Kahraman et al., 2021) 3D Solid Solve design problems incorporating deep learning Voxel grid Based on VAE & GAN
(Sebestyen et al., 2021) 3D Solid New way to design models Voxel grid VAE
(Guida, 2023) Text-3D Solid Multimodal generation with text input - Diffusion model
Table 3. An overview of 3D generative approaches for shape generation, each allowing the creation of editable models for explicit representations.
Method Names Publication & Year 3D Representation Deep Generative Model
3D GAN (Wu et al., 2016) NIPS 2016 Voxel grid GAN
Text2Shape (Chen et al., 2019) ACCV 2018 Voxel grid GAN
PLATONICGAN (Henzler et al., 2019) CVPR 2019 Voxel grid GAN
IG GAN  (Lunz et al., 2020) arXiv 2020 Voxel grid GAN
Achlioptas et al.  (Achlioptas et al., 2018) ICML 2018 Point cloud GAN
Shu et al.  (Shu et al., 2019) CVPR 2019 Point cloud GAN
Get3d  (Gao et al., 2022) NeurIPS 2022 Mesh GAN
IM-Net  (Chen and Zhang, 2019) CVPR 2019 Neural field GAN
Kleineberg et al.  (Kleineberg et al., 2020) arXiv 2020 Neural field GAN
Brock et al.  (Brock et al., 2016) arXiv 2016 Voxel grid VAE
Autosdf  (Mittal et al., 2022) CVPR 2022 Voxel gird VAE
Sagnet  (Wu et al., 2019) TOG 2019 Voxel gird VAE
Li et al.  (Li et al., 2020) AAAI 2020 Voxel gird VAE
Pq-net  (Wu et al., 2020) CVPR 2020 Voxel gird VAE
AdversarialAE  (Zamorski et al., 2020) CVPR 2020 Voxel gird VAE
Multi-Chart  (Ben-Hamu et al., 2018) TOG 2018 Mesh VAE
SDM-NET  (Gao et al., 2019) TOG 2019 Mesh VAE
Tm-net  (Gao et al., 2021) TOG 2021 Mesh VAE
Polygen  (Nash et al., 2020) ICML 2020 Mesh VAE
PointFLow  (Yang et al., 2019) CVPR 2019 Point cloud Normaliz. flow model
CLIP-Forge  (Sanghi et al., 2022) CVPR 2022 Voxel grid Normaliz. flow model
DreamFusion (Poole et al., 2022) arXiv 2022 Mesh Diffusion model
PVD  (Zhou et al., 2021) CVPR 2021 Hybrid: point cloud-voxel Diffusion model
Magic3D  (Lin et al., 2023) arXiv 2022 Neural field-Mesh Diffusion model
LION  (Zeng et al., 2022) arXiv 2022 Mesh Diffusion model
Point-E  (Nichol et al., 2022b) arXiv 2022 Neural field-Point cloud Diffusion model
Magic123 (Qian et al., 2023) arXiv 2023 Mesh Diffusion model
Prolificdreamer  (Wang et al., 2023) arXiv 2023 Mesh Diffusion model
Hifi-123 (Yu et al., 2024) arXiv 2024 Mesh Diffusion model
MVDream  (Shi et al., 2024) arXiv 2024 Mesh Diffusion model
Table 4. An overview of 3D generative approaches of 3D-aware image synthesis. ‘Single/multiple’ represents the result generated by a single image adopting a sample of single-view or multiple images adopting multiple-view images. ‘Geometry’ indicates whether this method allows the export to mesh. ‘Editability’ refers to the ability to directly manipulate or modify the synthesized output, such as adding or removing objects or adjusting their properties. ‘Controllability’ refers to the ability to adjust variables during the generation process, such as camera pose, object position, and lighting.
Method Names Publication & Year 3D Repre- sentation Single/multi- ple-view Geometry Editability Controllability Highlight
Camera Object
NeRF  (Mildenhall et al., 2021) CVPR 2019 Neural field multiple - - ✓- 3D Novel View Synthesis
HoloGAN  (Nguyen-Phuoc et al., 2019) CVPR 2019 Voxel grid single - - -
Pi-GAN  (Chan et al., 2021) CVPR 2021 Neural field single - - - -
Giraffe  (Niemeyer and Geiger, 2021) CVPR 2021 Neural field single - -
StyleSDF  (OrEl et al., 2022) CVPR 2022 Neural field single - - -
StyleNeRF  (Gu et al., 2021) arXiv 2021 Neural field single - - -
DreamField  (Jain et al., 2022) CVPR 2022 Neural field multiple - Text input (CLIP)
DreamFusion  (Poole et al., 2022) arXiv 2022 Neural field single Text input
CLIP-NeRF  (Wang et al., 2022) CVPR 2022 Neural field single - Text input (CLIP)
Table 5. The related works in architecture are classified into three categories: 2D to 3D transposition, 3D solid generation, and text-based 3D form generation. The ‘Category’ column denotes the generation approaches: ‘2D to 3D’ refers to generating 3D forms from 2D images; ‘3D Solid’ involves directly generating 3D forms using DGMs such as GANs, VAEs, 3D-aware image synthesis, and diffusion models; ‘Text-3D Solid’ indicates that the 3D form generation process incorporates text input to control the output. The ‘Objective’ column describes the research directions and goals. The ‘Methodology’ column outlines the proposed workflow for the generation process. ‘3D Representation’ indicates the formats of the generated data, including point cloud, voxel grid, and mesh. ‘Generative Model’ specifies the type of deep generative model used in each work.
(del Campo et al., 2019) 2D to 3D Test AI agency in design Utilizing Style Transfer to train two datasets Baroque and Modern images as a basis to form a 3D model - -
(Özel, 2020) 2D to 3D HCI in urban design Utilizing Style Transfer to generate different stylized images and generate 3D geometry through procedural modeling - -
(Ren and Zheng, 2020) 2D to 3D Test AI agency in design Utilizing Style Transfer to replace pixels with voxelization units to generate 3D forms - -
(Liu et al., 2021) 2D to 3D Toolkits for 3D generation Utilizing Style Transfer to assist the generation of 3D structure from 2D images - -
(Yu, 2020) 2D to 3D Generate building massing Utilizing Pix2Pix GAN to generate plan pattern and section pattern, then converted to 3D massing - -
(Di Carlo et al., 2022) 2D to 3D Generate building massing Utilizing Pix2Pix GAN to generate urban morphology to create building massing - -
(Zhang, 2019) 2D to 3D Form finding to assist design Slicing a 3D model and trained with different combinations of 2D StyleGANnetworks, and finally stitching into a 3D model - -
(Zhang, 2019) 2D to 3D Form finding to assist design 3D model generation based on 2D plan and section using Style Transfer - -
(Zhang and Huang, 2021) 2D to 3D Form finding to assist design Combining the spatial sequence information to generate 3D form from 2D images through multi-level deep generative networks such as StyleGAN - -
(Bank et al., 2022) 2D to 3D Human and neural network interface Utilizing 3D solid for training to map spatial semantics to a latent space assembled using point cloud representations Point cloud GAN
(Asmar and Sareen, 2020) 2D to 3D Integrate latent space in design GAN allows for navigation in the latent space to create digital designs using vector arithmetic and interpolation techniques, then converting resulting images to 3D voxel structures Voxel grid GAN
(Huang et al., 2021) 2D to 3D Recognize 2D pattern to 3D form Utilizing Latent space rotation and perspective projection to generate 3D model Voxel grid GAN
(Kim and Huang, 2022) 2D to 3D Recognize 2D pattern to 3D form Utilizing Latent space rotation and perspective projection to generate 3D model Voxel grid GAN
(Ondrej Vesel‘y, 2022) 2D to 3D Generate building massing Employing voxelization and vectorization techniques on images generated by the Pix2Pix GAN model to generate 3D building massing Voxel grid GAN
(Çakmak, 2022) 3D Solid Extend design cognition Utilizing a GAN with a pair of encoder-decoder to process datasets and generate new 3D models, resulting in different 3D representations like point cloud and mesh Point cloud/Mesh GAN (with Autoencoder)
(Mueller, 2023) 3D Solid Generate building massing Utilizing BuildingNet dataset (Selvaraju et al., 2021) as a basis to train urban morphology data to generate 3D building massing - GAN
(Miguel et al., 2019) 3D Solid Generation, manipulation and form finding of structural typologies Utilizing VAE to learn continuous latent space to generate new geometries Voxel grid VAE
(Kahraman et al., 2021) 3D Solid Solve design problems incorporating deep learning Develop post-integration in the generation workflow using VAE and GAN to solve complex multi-objective with abstract design criteria Voxel grid Base on VAE & GAN
(Sebestyen et al., 2021) 3D Solid New way to design models Utilizing VAE to encode and decode geometries through dimensionality manipulation Voxel grid VAE
(Liu et al., 2024) 3D Solid Form finding to assist in design Developing a tool with a user interface (UI) to generate 3D point clouds using a diffusion model. Point cloud Diffusion model
(Sebestyen et al., 2023) 3D Solid Form finding to assist design Utilizing diffusion model to generate 3D forms with voxel grids through parametric design as dataset Voxel grid Diffusion model
(Guida et al., 2023) Text-3D Solid Multimodal generation with text input Utilizing diffusion model and NeRF to generate 3D forms from text input with multimodality Mesh Diffusion model & NeRF
(Koh, 2023) Text-3D Solid Test the connection between text and 3D form Employing a specialized prompt to manipulate DreamField to generate multiple design alternatives for architects Mesh Diffusion model
(Guida, 2023) Text-3D Solid Multimodal generation with text input Utilizing diffusion model to generate 3D forms from text input with multimodality - Diffusion model